OpenStudy (anonymous):

The mean of five observation is 17. If the mean of first three observation is 15 and the last three is 18. Find the third observation.

OpenStudy (anonymous):

See, Sum of Observation = Mean*Number of Observation..

OpenStudy (anonymous):

i did'nt understand the question! please can u explain.

OpenStudy (anonymous):

Sum of 5 observation = 17*5 = 85 first(2) + x = 15*3 = 45 (x is the third observation) ---------1 x + next(2) = 18*3 = 54 -------------2 Now add both the equations and subtract x: (first(2) + x) + (x + next(2)) - x = 85 45 + 5 4 - x = 85 99 - x = 85 x = 14 So, the third observation is 14...

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

Yes why not.

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

there is one more to solve, can u help?

OpenStudy (anonymous):

Yes sure..

OpenStudy (anonymous):

The mean of ten observations is found to be 10. Later on, it was found that the observation 21was misread as 12. Find the correct mean.

OpenStudy (anonymous):

See first find the sum of observation.. You have 10 observation and mean = 10 use the above formula..

OpenStudy (anonymous):

sum of observation is 100.

OpenStudy (anonymous):

In sum there is one term that is 12 and if we want to make it 21 so what number i will add to 12 to get 21?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

after that?

OpenStudy (anonymous):

hey! please help me!!!!!!

OpenStudy (anonymous):

Yes add 9 to the sum of observation..

OpenStudy (anonymous):

109

OpenStudy (anonymous):

109-100=9 like that??

OpenStudy (anonymous):

109 is the right one..

OpenStudy (anonymous):

answer =109

OpenStudy (anonymous):

yaa?

OpenStudy (anonymous):

now divide it by 10 to get the mean..

OpenStudy (anonymous):

10.9

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

Yes this is it..

OpenStudy (anonymous):

thnx :)